Prove that a sequence that diverges to +∞ (resp. −∞) is divergent according to the previous definition


1
Expert's answer
2021-02-02T01:38:21-0500

prev definition is not mentioned.

However trying to make sense from the question asked, it was figured, it is correctly formulated as a sequence converges to infinity. Then this sequence diverges i.e. doesn't converge.

So a sequence converges to infinity means given "M>0," there exists a natural number "N"

such that "\forall n\geq N, x_n>M." Here the sequence is denoted as "\{x_n\}." Now this sequence cannot converge to any point since if it converges to "x," then we take "\epsilon =1." So for some natural "N," "n\geq N\Rightarrow |x_n-x|<1." Hence we get, "x_n<x+1." This contradicts the previous definition by taking, "M=x+1." Hence the sequence diverges according to the definition that diverges means not convergent.